Zeronknight (zeronknight):

Simplify the following ratio. \frac{70}{105} A. \frac{7}{10} B. \frac{2}{3} C. \frac{10}{15} D. \frac{14}{21}

Zeronknight (zeronknight):

\[\frac{70}{105}\]

Zeronknight (zeronknight):

\[A. \frac{7}{10}\]

Zeronknight (zeronknight):

\[B. \frac{2}{3}\]

Zeronknight (zeronknight):

\[C. \frac{10}{15}\]

Zeronknight (zeronknight):

\[D. \frac{14}{21}\]

OpenStudy (anonymous):

well, find a number that both 70 and 105 can be divided by